How many milliliters of distilled water must be added to 50 milliliters of 40% acid solution to obtain 20% acid solution?
x(0)+ 50(.4)=x+50(.2)
x+20=x+10

Answer by mananth(16946)

--- percent ---------------- Amount
Solution 40 ---------------- 50 ml
Water 0 ---------------- x ml
Mixture 20 ---------------- 50 + x ml

40 * 50 + 0 x = 20 ( 50 + x )
2000 + 0 x = 1000 + 20 x
0 x -20 x = -2000 + 1000
-20 x = -1000
/ -20
x = 50 ml Water

m.ananth@hotmail.ca